# Comment on Marek Czachor article entitled “On Relativity of Quantumness as Implied by Relativity of Arithmetic and Probability” beginequation0.25em]

### v1

@@ Abstract
Czachor’s model of hierarchical arithmetics begins with a valid formal premise but fixes the key probability mapping $g$ by importing the Born rule and Fubini–Study metric from standard quantum mechanics, where Born probabilities are Kolmogorov within a fixed measurement context. This $g$ is then applied in a non-Newtonian hidden-variable setting, producing a hybrid framework whose agreement with quantum correlations is built in by design, not derived from new physics, and thus does not constitute a genuine counterexample to Bell’s theorem.
@@
Lemma 1. The map should satisfy the natural symmetry g(p)+g(1p)=1g(p)+g(1-p)=1.
@@
Here $\theta$ is the angle between the corresponding Bloch vectors of the pure states $|a\rangle$ and $|b\rangle$. Caveat. The FS formula above applies to pure states. For mixed states the natural geometric replacement involves quantum fidelity (Bures angle), and the calibration generalizes accordingly. In particular, the Bures angle is $\mathcal{B}(\rho,\sigma)=\arccos\!\sqrt{F(\rho,\sigma)}$ with $F$ the Uhlmann fidelity.
@@
Remark. The linear relation $\theta=\pi(1-p)$ is a modeling choice (simple and symmetric), not a consequence of Lemma 1. Many other choices would lead to different $g$’s. (Indeed, any monotone $\theta:[0,1]\to[0,\pi]$ with $\theta(0)=\pi,\ \theta(1)=0$ yields an admissible $g$ via $g(p)=\cos^{2}\!\big(\tfrac{\theta(p)}{2}\big)$.)
@@
For $g_{1}$, one checks
$g_{1}(p)+g_{1}(1-p)=3p^{2}-2p^{3}+3(1-p)^{2}-2(1-p)^{3}=1.$
Monotonicity. Because $s^{\prime}(p)=\tfrac{\pi}{2}\sin(\pi p)>0$ on $(0,1)$ and $\sin^{2}x$ is strictly increasing on $0,\tfrac{\pi}{2}]$, $g_{\text{alt}}$ is strictly increasing on $[0,1]$ and hits the boundaries. Monotonicity of $g_{1}$. Since $g_{1}^{\prime}(p)=6p-6p^{2}=6p(1-p)\geq 0$ on $[0,1]$ (strictly >0>0 on $(0,1)$), $g_{1}$ is strictly increasing and hits the boundaries $g_{1}(0)=0$, $g_{1}(1)=1$. Both examples satisfy g(p)+g(1p)=1g(p)+g(1-p)=1 and reach g(0)=0,g(1)=1g(0)=0,\ g(1)=1; they genuinely differ from $g(p)=\sin^{2}(\tfrac{\pi p}{2})$. (Note that $\tfrac{1-\cos(\pi p)}{2}\equiv\sin^{2}(\tfrac{\pi p}{2})$, so that cosine form is not independent.)
Figure 1: Example mappings $g(p)$ satisfying Lemma 1 constraints. Czachor’s choice $g(p)=\sin^{2}\!\big(\frac{\pi p}{2}\big)$ (solid) is one among many admissible forms, including the polynomial $3p^{2}-2p^{3}$ (dashed) and a distinct, symmetric sine-embedded variant $g_{\text{alt}}(p)=\sin^{2}\!\big(\frac{\pi}{4}[1+\sin(\pi(p-\frac{1}{2}))]\big)$ (dotted). All three are continuous, strictly increasing, satisfy g(p)+g(1p)=1g(p)+g(1-p)=1, and interpolate $(0,0)$ to $(1,1)$.
@@
On the hidden-variable side we now operate with a non-Kolmogorov probability calculus, so standard Bell-inequality derivations don’t go through as usual. [^1] Equivalent formulations of Bell’s premises include factorizability or outcome & parameter independence. We use the common shorthand of locality, realism, and independence of settings. On the observable side we have chosen $g$ to match quantum statistics.
@@
If one truly abandons observable-level Kolmogorov structure, there is no special reason to prefer $g(p)=\sin^{2}(\tfrac{\pi p}{2})$; Lemma 1 leaves infinitely many choices. The close fit to quantum correlations is therefore largely by construction, not a new physical derivation.
In that sense, the proposal is not a counterexample to Bell’s theorem in its original scope. It changes a premise (the probability calculus) and then reuses a quantum-calibrated $g$, so agreement with QM is built in rather than emergent.
@@ Conclusion
Czachor’s proposal starts from a formally coherent observation about isomorphic arithmetic hierarchies, but the specific framework he constructs does not hold up under closer scrutiny. Lemma 1, while central to the argument, leaves the mapping $g$ largely undefined—yet Czachor proceeds to anchor it using the Born rule and the Fubini–Study metric, both of which originate in standard quantum theory where Born probabilities are Kolmogorov within a fixed measurement context. This step, crucially, reintroduces the very arithmetic formalism that the hierarchy model was intended to move beyond. What emerges is not a clean break from classical probability, but rather a hybrid Kolmogorov/non-Kolmogorov model—an uneasy blend wherein the core mechanism relies on principles it ostensibly rejects.
Within this setup, the emergence of quantum singlet correlations isn’t a surprising result, but an engineered outcome—an artefact of how $g$ is chosen. Because this choice is made rather than derived, the model fails to qualify as a counterexample to Bell’s theorem. None of Bell’s foundational premises—locality, realism, or the independence of measurement settings—are actually negated or circumvented. Instead, the approach modifies the arithmetic scaffold and surreptitiously reinstates standard quantum mechanical constructs. (In one precise formulation, Bell’s theorem follows from classical Kolmogorov probability with factorizability and measurement independence.) This doesn’t challenge Bell’s theorem so much as reframe probability theory itself. The apparent reproduction of quantum correlations only shows that, by redefining the underlying calculus, one can reassemble results that Bell showed to be unattainable within classical constraints. In that regard, Czachor’s construction serves more as a reinterpretation of how probabilities are modeled than a disproof of Bell’s logic.
@@
“…I am changing the paradigm of physics. I propose that, just as Einstein incorporated geometry into physics so that geometry ceased to be something abstract, Kantian and a priori, and instead became a branch of physics — which is the essence of general relativity — at this moment I claim that an analogous status applies to arithmetic structures. They should not be treated as something given a priori, but as something to be used in the most general way possible, and then let experiment decide what is concretely valid. In this sense it is a new paradigm, and theorists of science say that a revolution takes place when the paradigm changes. Such a paradigm shift was relativity theory …I am incorporating arithmetic into physics …perhaps nature adds and multiplies differently than we imagine …”
The analogy to Einstein’s revolution is certainly evocative—but it oversells the current state of the formalism. Our examination of Lemma 1 in [1] reveals that $g(p)$ is far from uniquely determined. An infinite family of admissible functions fits the loose constraints, and Czachor’s sinusoidal choice is lifted wholesale from conventional quantum mechanics. The framework thereby ends up recycling elements it initially set out to replace.
A parallel issue appears across Czachor’s broader body of work on generalized calculus [2, 3], where different bijections $f$ can produce either physically consistent outcomes (such as valid relativistic velocity addition via $f(\beta)=\mathrm{arctanh}(\beta)$) or unphysical results (e.g. $f(\beta)=\beta^{3}$ yielding superluminal speeds). For instance, under the induced operation $\beta_{1}\oplus\beta_{2}=f^{-1}(f(\beta_{1})+f(\beta_{2}))$, taking $\beta_{1}=\beta_{2}=0.9$ with $f(\beta)=\beta^{3}$ gives $f(\beta_{i})=0.729$ and $f(\beta_{1})+f(\beta_{2})=1.458$, hence $\beta_{1}\oplus\beta_{2}=(1.458)^{1/3}\approx 1.127>1$ (superluminal). The fact that both physically acceptable and unacceptable cases emerge from the same formal approach highlights the absence of internal selection criteria—a significant limitation when claiming to offer a foundational overhaul of physics.

### v2

@@ Abstract
Czachor’s model of hierarchical arithmetics begins with a valid formal premise but fixes the key probability mapping $g$ by importing the Born rule and Fubini–Study metric from standard quantum mechanics, where Born probabilities are Kolmogorov within a fixed measurement context. This $g$ is then applied in a non-Newtonian hidden-variable setting, producing a hybrid framework whose agreement with quantum correlations is built in by design, not derived from new physics, and thus does not constitute a genuine counterexample to Bell’s theorem. Moreover, the construction changes a standard premise used in one precise formulation of Bell’s theorem (classical Kolmogorov probability), and therefore lies outside that theorem’s original scope.
@@
Assumption 1 (Complement preservation). We explicitly assume the regraduation respects complements,
g(p)+g(1p)=1.g(p)+g(1-p)=1.
Immediate consequence. By Assumption 1, $g(\tfrac{1}{2})=1-g(\tfrac{1}{2})$, hence $g(\tfrac{1}{2})=\tfrac{1}{2}$.
@@
Here $\theta$ is the angle between the corresponding Bloch vectors of the pure states $|a\rangle$ and $|b\rangle$.
Caveat. The FS formula above applies to pure states. For mixed states a natural geometric replacement uses quantum fidelity (Bures angle) $\mathcal{B}(\rho,\sigma)=\arccos\!\sqrt{F(\rho,\sigma)}$. However, there is no universal identity of the form $P=\cos^{2}\mathcal{B}$ for general mixed states; probabilities remain $P(b|\rho)=\mathrm{tr}[\rho\,\Pi_{b}]$ for a specified POVM, so any “calibration” beyond the pure-state case must state the measurement model explicitly.
@@
Remark. The linear relation $\theta=\pi(1-p)$ is an illustrative modeling choice, not a consequence of Assumption 1. Many other choices would lead to different $g$’s. However, to satisfy Assumption 1 (complement preservation) the parametrization must obey the symmetry
$\theta(1-p)=\pi-\theta(p),$
since then
$g(p)+g(1-p)=\cos^{2}\!\frac{\theta(p)}{2}+\cos^{2}\!\frac{\theta(1-p)}{2}=\cos^{2}x+\sin^{2}x=1\quad\text{with }x=\tfrac{\theta(p)}{2}.$
Equivalently, one may define $g$ arbitrarily on $0,\tfrac{1}{2}]$ (subject to continuity/monotonicity, $g(0)=0$, and $g(\tfrac{1}{2})=\tfrac{1}{2}$ to ensure continuity at the join) and extend by g(1p)=1g(p)g(1-p)=1-g(p). Thus not every monotone $\theta$ yields an admissible $g$; the symmetry (or complement-extension) is required.
@@
Monotonicity. Because $s^{\prime}(p)=\tfrac{\pi}{2}\sin(\pi p)>0$ on $(0,1)$, $s([0,1])=[0,1]$, and $\sin^{2}x$ is strictly increasing on $(0,\tfrac{\pi}{2})$, the inner argument $\tfrac{\pi}{2}s(p)\in[0,\tfrac{\pi}{2}]$ is non-constant, so $g_{\text{alt}}$ is strictly increasing on $[0,1]$ and hits the boundaries. Monotonicity of $g_{1}$. Since $g_{1}^{\prime}(p)=6p-6p^{2}=6p(1-p)\geq 0$ on $[0,1]$ (strictly >0>0 on $(0,1)$), $g_{1}$ is strictly increasing and hits the boundaries $g_{1}(0)=0$, $g_{1}(1)=1$. Both examples satisfy Assumption 1 and reach g(0)=0,g(1)=1g(0)=0,\ g(1)=1; they genuinely differ from $g(p)=\sin^{2}(\tfrac{\pi p}{2})$. (Note that $\tfrac{1-\cos(\pi p)}{2}\equiv\sin^{2}(\tfrac{\pi p}{2})$, so that cosine form is not independent.)
Figure 1: Example mappings $g(p)$ satisfying Assumption 1 (complement preservation), continuity, and strict monotonicity. Czachor’s choice $g(p)=\sin^{2}\!\big(\frac{\pi p}{2}\big)$ (solid) is one among many admissible forms, including the polynomial $3p^{2}-2p^{3}$ (dashed) and a distinct, symmetric sine-embedded variant $g_{\text{alt}}(p)=\sin^{2}\!\big(\frac{\pi}{4}[1+\sin(\pi(p-\frac{1}{2}))]\big)$ (dotted). All three interpolate $(0,0)$ to $(1,1)$.
@@
On the hidden-variable side we now operate with a non-Kolmogorov probability calculus, so standard Bell-inequality derivations don’t go through as usual. [^1] Equivalent formulations of Bell’s locality condition include factorizability, i.e. the conjunction of outcome independence and parameter independence, within the classical (Kolmogorov) framework. We use the common shorthand of locality, realism, and independence of settings. On the observable side we have chosen $g$ to match quantum statistics.
@@
If one truly abandons observable-level Kolmogorov structure, there is no special reason to prefer $g(p)=\sin^{2}(\tfrac{\pi p}{2})$; Assumption 1 leaves infinitely many choices. The close fit to quantum correlations is therefore largely by construction, not a new physical derivation.
In that sense, the proposal is not a counterexample to Bell’s theorem in its original scope. It modifies a standard premise (classical Kolmogorov probability) and then reuses a quantum-calibrated $g$, so agreement with QM is built in rather than emergent.
@@ Conclusion
Czachor’s proposal starts from a formally coherent observation about isomorphic arithmetic hierarchies, but the specific framework he constructs does not hold up under closer scrutiny. Assumption 1, while central to the argument, leaves the mapping $g$ largely underdetermined—yet Czachor proceeds to anchor it using the Born rule and the Fubini–Study metric, both of which originate in standard quantum theory where Born probabilities are Kolmogorov within a fixed measurement context. This step, crucially, reintroduces the very arithmetic formalism that the hierarchy model was intended to move beyond. What emerges is not a clean break from classical probability, but rather a hybrid Kolmogorov/non-Kolmogorov model—an uneasy blend wherein the core mechanism relies on principles it ostensibly rejects.
Within this setup, the emergence of quantum singlet correlations isn’t a surprising result, but an engineered outcome—an artefact of how $g$ is chosen. Because this choice is made rather than derived, the model fails to qualify as a counterexample to Bell’s theorem. Since it modifies a standard probability premise in the theorem’s usual (Kolmogorov-based) derivations, it falls outside that scope; the approach reframes the premises rather than refuting them. The apparent reproduction of quantum correlations only shows that, by redefining the underlying calculus, one can reassemble results that Bell showed to be unattainable within classical constraints. In that regard, Czachor’s construction serves more as a reinterpretation of how probabilities are modeled than a disproof of Bell’s logic.
@@
“…I am changing the paradigm of physics. I propose that, just as Einstein incorporated geometry into physics so that geometry ceased to be something abstract, Kantian and a priori, and instead became a branch of physics — which is the essence of general relativity — at this moment I claim that an analogous status applies to arithmetic structures. They should not be treated as something given a priori, but as something to be used in the most general way possible, and then let experiment decide what is concretely valid. In this sense it is a new paradigm, and theorists of science say that a revolution takes place when the paradigm changes. Such a paradigm shift was relativity theory …I am incorporating arithmetic into physics …perhaps nature adds and multiplies differently than we imagine …"
The analogy to Einstein’s revolution is certainly evocative—but it oversells the current state of the formalism. Our examination of Assumption 1 in [1] reveals that $g(p)$ is far from uniquely determined. An infinite family of admissible functions fits the loose constraints, and Czachor’s sinusoidal choice is lifted wholesale from conventional quantum mechanics. The framework thereby ends up recycling elements it initially set out to replace.
A parallel issue appears across Czachor’s broader body of work on generalized calculus [2, 3], where different bijections $f$ can produce either physically consistent outcomes (such as valid relativistic velocity addition via $f(\beta)=\operatorname{artanh}(\beta)$) or unphysical results. In general, the induced operation $\beta_{1}\oplus\beta_{2}=f^{-1}\!\big(f(\beta_{1})+f(\beta_{2})\big)$ is only a partial operation on (1,1)(-1,1) unless $\operatorname{Im}(f)$ is closed under addition (e.g., $f=\operatorname{artanh}$ has $\operatorname{Im}(f)=\mathbb{R}$). For instance, taking $\beta_{1}=\beta_{2}=0.9$ with $f(\beta)=\beta^{3}$ gives $f(\beta_{i})=0.729$ and $f(\beta_{1})+f(\beta_{2})=1.458$, which lies outside $\operatorname{Im}(f)=(-1,1)$. Hence $\beta_{1}\oplus\beta_{2}$ is not defined without extending $f^{-1}$ beyond its natural domain. If one does extend $f^{-1}$ to $\mathbb{R}$, one obtains $(1.458)^{1/3}\approx 1.134$, which merely signals loss of closure of the operation rather than a physical “superluminal sum.” The fact that both physically acceptable and unacceptable cases emerge from the same formal approach highlights the absence of internal selection criteria—a significant limitation when claiming to offer a foundational overhaul of physics.